OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). a.)-40x + 24 b.)-40x + 31 c.)-40x + 64 d.)-40x + 70

OpenStudy (anonymous):

f(x) = -4x + 7 g(x) = 10x - 6 f(g(x))=-4(g(x) +7= -4(10x-6)+7

OpenStudy (anonymous):

f(g(x))=-4(g(x) +7= -4(10x-6)+7= -40 x + 24 +7=-40 x+ 31

OpenStudy (anonymous):

ok thanks @eliassaab

OpenStudy (anonymous):

yw